OpenStudy (anonymous):

hey there can somebody help me to find the x-intercept of this equation f(x)=(x+5)-4

OpenStudy (anonymous):

y = (x+5)-4 x intercept is when y=0 0 = (x+5)-4 4= (x+5) 4 - 5 = x x = -1 therefore, the x intercept would be (-1,0)

OpenStudy (lalaly):

set f(x)=0 (x+5)-4=0 x+1=0 x=-1 so the x intercept is (-1,0)

OpenStudy (anonymous):

-1